OpenStudy (anonymous):

What is the factor of x^3-4x^2+x+6?

OpenStudy (anonymous):

here is a good gimmick to know: it is going to be easy to factor if you can find the zeros math teachers usually make a zero of 1 or -1 because those are the easiest, so check those first

OpenStudy (anonymous):

1 does not work, because if you plug it in to \(x^3-4x^2+x+6\) you get \[1-4+1+6\] which is not zero

OpenStudy (anonymous):

but \(-1\) does work, you get \[-1-4-1+6=0\]

OpenStudy (anonymous):

therefore you know it factors \[x^3-4x^2+x+6=(x+1)(something)\] and you can find the "something" either by division (synthetic division is easiest) or by thinking

OpenStudy (anonymous):

your choice

OpenStudy (anonymous):

What divisor can i use in synthetic division?

OpenStudy (anonymous):

you want to factor \(x^3-4x^2+x+6\) as \((x+1)(whatever)\) this makes \[whatever=(x^3-4x^2+x+6)\div (x+1)\]

OpenStudy (anonymous):

you can also just figure out what it has to be in your head more or less

OpenStudy (anonymous):

i got 3x

OpenStudy (anonymous):

how? \[x^3-4x^2+x+6=3x(x-1)?\] that can't be correct, can it?

OpenStudy (anonymous):

No, it's wrong. :3

OpenStudy (anonymous):

actually i meant to write \[x^3-4x^2+x+6=3x(x+1)?\] but that is not correct either your "whatever" needs to be a polynomial of degree 2

OpenStudy (anonymous):

it is clear what you have to do? you have to write \[x^3-4x^2+x+6=(x+1)(ax^2+bx+c)\]